Consider the following regression equation Ŷ=18.92+1.11X The residual when X = 40 and the Y = 60 is?


1
Expert's answer
2021-08-11T11:40:20-0400

Solution:

Ŷ=18.92+1.11X

Put X=40

Ŷ=18.92+1.11(40) = 63.32

Given, Y = 60

So, residual = Ŷ-Y = 63.32 - 60 = 3.32